# bibliography.bib

### v1

@@ Abstract
Humans have exceptional tactile sensing capabilities, which they can leverage to solve challenging, partially observable tasks that cannot be solved from visual observation alone. Research in tactile sensing attempts to unlock this new input modality for robots. Lately, these sensors have become cheaper and, thus, widely available. At the same time, the question of how to integrate them into control loops is still an active area of research, with central challenges being partial observability and the contact-rich nature of manipulation tasks. In this study, we propose to use Reinforcement Learning to learn an end-to-end policy, mapping directly from tactile sensor readings to actions. Specifically, we use Dreamer-v3 on a challenging, partially observable robotic insertion task with a Franka Research 3, both in simulation and on a real system. For the real setup, we built a robotic platform capable of resetting itself fully autonomously, allowing for extensive training runs without human supervision. Our preliminary results indicate that Dreamer is capable of utilizing tactile inputs to solve robotic manipulation tasks in simulation and reality. Furthermore, we find that providing the robot with tactile feedback generally improves task performance, though, in our setup, we do not yet include other sensing modalities. In the future, we plan to utilize our platform to evaluate a wide range of other Reinforcement Learning algorithms on tactile tasks.
@@ I Introduction
A central challenge of many manipulation tasks is partial observability, as object tracking is often inaccurate, especially due to occlusion caused by the robot’s end-effector. While tactile sensors can provide crucial information about the points of contact, the information obtained from them will usually still be incomplete, rendering the decision-making problem a Partially Observable Markov Decision Process (POMDP). Yet, reinforcement learning (RL) algorithms commonly rely on the assumption that the state-space is Markovian and fully observable [schulman2017ppo, haarnoja2018sac, fujimoto2018td3, chen2021redq, bhatt2023crossq]. Recently, a number of RL algorithms have been proposed that tackle POMDPs by learning a latent state-space model of the environment [Hafner_2019_LatDynPixels, lee2020stochastic, Havens_Ouyang_LearningLatentStateSpaces, Rybkin_2021_MBRLViaLatentSpaceColloc, Nguyen_2021_TemporalPredictiveCodingForMB, Gelada_et_al_2019, Hafner_2020_DreamToControl, Hafner_2020_MasteringAtari, Deng_2021_DreamerPro, Wu_hafner_2022_DayDreamer].
@@ II Robotic Platform
The robotic setup, consisting of a 7 DoF Franka Research 3, a circular peg, and a base plate with an insertion hole, is presented in Fig. 1. We mount a vision-based GelSight Mini tactile sensor [yuan2017gelsight] sensor in the fingertips of the robot’s gripper. A key challenge is the development of a reset procedure that places the agent back in its initial state, with the peg properly placed in its gripper. To prevent the loss of the cylinder if the peg slips from the gripper, we suspend the peg with a thin thread connected to the gripper. Upon losing contact, the peg hangs from the thread. The robot then moves to a reset box, lowers the peg inside, re-grasps, and returns to its initial position above the insertion hole. This resetting procedure enables the agent to train consecutively for multiple days without requiring human intervention.
The agent’s observation comprises a RGB image from the Gelsight sensor and the end-effector position in Cartesian coordinates at . The policy returns the relative new positions of the end-effector at . These target positions are then fed into the franky [franky] control library, which internally uses ruckig [berscheid2021jerk] for smooth motion planning and Franka’s internal cartesian impedance controller for execution. We ensure safe exploration of the environment by restricting the workspace to be in the vicinity of the hole. For continuous tracking of the cylinder pose, we utilize OptiTrack [optitrack]. This information is only used for evaluation of the robot’s performance and not as an input to the policy.
@@
As shown in Table I, the reward comprises of four components: (i) proximity to the goal, (ii) a terminal reward upon reaching the goal , (iii) a terminal penalty for leaving the workspace, and (iv) an action penalty to encourage smooth motions. The policy training includes several neural network models: encoder and decoder networks that project between the observation space and the 4160-dimensional latent space, a recurrent state-space model, a reward model in the latent space, and the policy.
In addition, we develop a digital twin of our setup in PyBullet [coumans2016pybullet], tactile sensors are simulated with Taxim [taxim].
@@ III Empirical Evaluations
We train Dreamer from scratch, both in simulation and reality. In either environment, we conduct experiments with and without tactile sensors to evaluate whether tactile sensing yields a benefit for our task.
@@ IV Conclusion & Future Work
In this work, we developed a robot platform that facilitates autonomous training of an RL agent on a Franka Research 3 robot to investigate whether reinforcement learning is a viable option for incorporating tactile feedback in control loops. To ensure reliable and safe training, we implement a robust resetting routine and restrict the action space, enabling autonomous learning of an insertion policy without human supervision. Leveraging tactile feedback from a vision-based tactile sensor attached to the gripper, our policy autonomously learns to insert a peg into a hole using the model-based RL approach Dreamer. Our simulation results suggest that the inclusion of tactile information significantly enhances learning outcomes. With our platform, we intend to benchmark other RL algorithms on tactile insertion tasks to get a clear understanding of which methods are capable of dealing with the inherent complexity of tactile manipulation tasks in the real world. Furthermore, we plan to increase task complexity in future experiments, enabling a thorough exploration of touch’s role in dexterous manipulation.

### v2

@@ Abstract
Humans have exceptional tactile sensing capabilities, which they can leverage to solve challenging, partially observable tasks that cannot be solved from visual observation alone. Research in tactile sensing attempts to unlock this new input modality for robots. Lately, these sensors have become cheaper and, thus, widely available. But, how to integrate them into control loops is still an active area of research, with central challenges being partial observability and the contact-rich nature of manipulation tasks. In this study, we propose to use Reinforcement Learning to learn an end-to-end policy, mapping directly from tactile sensor readings to actions. Specifically, we use Dreamer-v3 on a challenging, partially observable robotic insertion task with a Franka Research 3, both in simulation and on a real system. For the real setup, we built a robotic platform capable of resetting itself fully autonomously, allowing for extensive training runs without human supervision. Our initial results show that Dreamer is capable of utilizing tactile inputs to solve robotic manipulation tasks in simulation and reality. Further, we find that providing the robot with tactile feedback generally improves task performance, though, in our setup, we do not yet include other sensing modalities. In the future, we plan to utilize our platform to evaluate a wide range of other Reinforcement Learning algorithms on tactile tasks.
@@ I Introduction
A central challenge of many manipulation tasks is partial observability, as object tracking is often inaccurate, especially due to occlusion caused by the robot’s end-effector. While tactile sensors can provide crucial information about the points of contact, the information obtained from them will usually still be incomplete, rendering the decision-making problem a Partially Observable Markov Decision Process (POMDP). Yet, reinforcement learning (RL) algorithms and model-based RL algorithms commonly rely on the assumption that the state-space is Markovian and fully observable [schulman2017ppo, haarnoja2018sac, fujimoto2018td3, schneider2022active, palenicek2023diminishing, bhatt2023crossq]. Recently, a number of RL algorithms have been proposed that tackle POMDPs by learning a latent state-space model of the environment [Hafner_2019_LatDynPixels, lee2020stochastic, Havens_Ouyang_LearningLatentStateSpaces, Rybkin_2021_MBRLViaLatentSpaceColloc, Nguyen_2021_TemporalPredictiveCodingForMB, Gelada_et_al_2019, Hafner_2020_DreamToControl, Hafner_2020_MasteringAtari, Wu_hafner_2022_DayDreamer].
@@ II Robotic Platform
The robotic setup, consisting of a 7 DoF Franka Research 3, a circular peg, and a base plate with an insertion hole, is presented in 1. We mount a vision-based GelSight Mini tactile sensor [yuan2017gelsight] sensor in the fingertips of the robot’s gripper. A key challenge is the development of a reset procedure that places the agent back in its initial state, with the peg properly placed in its gripper. To prevent the loss of the cylinder if the peg slips from the gripper, we suspend the peg with a thin thread connected to the gripper. Upon losing contact, the peg hangs from the thread. The robot then moves to a reset box, lowers the peg inside, re-grasps, and returns to its initial position above the insertion hole. This resetting procedure enables the agent to train consecutively for multiple days without requiring human intervention.
The agent’s observation comprises a RGB image from the Gelsight sensor and the end-effector position in Cartesian coordinates at . The policy returns the relative new positions of the end-effector at . These target positions are then fed into the franky [franky] control library, which internally uses ruckig [berscheid2021jerk] for smooth motion planning and Franka’s internal cartesian impedance controller for execution. We ensure safe exploration of the environment by restricting the workspace to be in the vicinity of the hole. For continuous tracking of the cylinder pose, we utilize OptiTrack. This information is only used for evaluation of the robot’s performance and not as an input to the policy.
@@
As shown in Table I, the reward comprises of four components: (i) proximity to the goal, (ii) a terminal reward upon reaching the goal , (iii) a terminal penalty for leaving the workspace, and (iv) an action penalty to encourage smooth motions. The policy training includes several neural network models: encoder and decoder networks that project between the observation space and the 4160-dimensional latent space, a recurrent state-space model, a reward model in the latent space, and the policy. In addition, we develop a digital twin of our setup in PyBullet [coumans2016pybullet], tactile sensors are simulated with Taxim [taxim].
@@ III Empirical Evaluations
We train Dreamer from scratch, both in simulation and reality. We conduct experiments with and without tactile sensors to evaluate the benefits of tactile sensing.
@@ IV Conclusion & Future Work
In this work, we developed a robot platform that facilitates autonomous training of an RL agent on a Franka Research 3 robot to investigate whether RL is a viable option for incorporating tactile feedback in control loops. To ensure reliable and safe training, we implement a robust resetting routine and restrict the action space, enabling autonomous learning of an insertion policy without human supervision. Leveraging tactile feedback from a vision-based tactile sensor attached to the gripper, our policy autonomously learns to insert a peg into a hole using the model-based RL approach Dreamer. Our simulation results suggest that the inclusion of tactile information significantly enhances learning outcomes. With our platform, we intend to benchmark other RL algorithms on tactile insertion tasks to get a clear understanding of which methods are capable of dealing with the inherent complexity of tactile manipulation tasks in the real world. Furthermore, we plan to increase task complexity in future experiments, enabling a thorough exploration of touch’s role in dexterous manipulation.